Traffic reduces Jim’s usual speed by 10mph, which lengthened his 200mph trip by 1 hour. Find his average speed.
Answer by Alan3354(69443)

Traffic reduces Jim’s usual speed by 10mph, which lengthened his 200mph trip by 1 hour. Find his average speed.
--------------
t = usual travel time
---
d = r*t = (r-10)*(t+1) = 200
r = 200/t
(200/t - 10)(t+1) = 200
200 + 200/t - 10t - 10 = 200
-10t - 10 + 200/t = 0
10t^2 + 10t - 200 = 0
t^2 + t - 20 = 0
(t+5)*(t-4) = 0
t = 4 hours
r = 200/4 = 50 mi/hr